For the function f(x) represented as y=x^2-4.
A) How do I write the equation defining the inverse of f(x)
B) How do I sketch the graphs of f and its inverse relation on the same axes?
C) How do I state whether the inverse relation is a function?
I am working on this at this moment and help is highly appreciated.
Thank You

Answer by stanbon(75887)

For the function f(x) represented as y=x^2-4.
A) How do I write the equation defining the inverse of f(x)
1st: Interchange x and y to get:
x = y^2-4
2nd: Solve for "y":
y^2 = x+4
y = +-sqrt(x+4)
-------------------------------
B) How do I sketch the graphs of f and its inverse relation on the same axes?

------------------------------------------------------
C) How do I state whether the inverse relation is a function?
It is not because there are x values in the domain of f^-1(x) that
have 2 associated y-values.
=================
Cheers,
Stan H.